### v2

# Diffusion in SPAD Signals

Abstract
We derive the likelihood of a raw signal in a single photon avalanche diode (SPAD), given a fixed photon flux. The raw signal comprises timing of detection events, which are nonlinearly related to the flux. Moreover, they are naturally stochastic. We then derive a score function of the signal. This is a key for solving inverse problems based on SPAD signals. We focus on deriving solutions involving a diffusion model, to express image priors. We demonstrate the effect of low or high photon counts, and the consequence of exploiting timing of detection events.
Index Terms:
Computational Photography, Inverse Problems, Diffusion, SPAD Models
1 Introduction
Single-Photon Avalanche Diode (SPAD) arrays offer capabilities that extend well beyond conventional sensors. Unlike standard detectors that accumulate charge, SPADs operate in Geiger-mode, enabling the detection of individual photon arrival events with picosecond-level timing precision and negligible read noise [1]. This architecture benefits fields requiring extreme sensitivity and/or high temporal resolution, e.g., in fluorescence lifetime imaging microscopy (FLIM), LiDAR, and non-line-of-sight imaging [2, 3].
One advantage of SPAD technology is its performance in photon-starved environments [4]. The ability to count single photons allows for imaging under extreme low-light conditions. It also supports ultra-high temporal rates, useful for capturing transient phenomena. However, these scanrios introduce significant reconstruction challenges. To leverage the speed of SPADs, exposure times are often reduced to the microsecond scale. There, the photon count per pixel is low [5].
Recovering high-fidelity intensity maps from such sparse, noisy, and discrete measurements is a challenge. Recently, diffusion probabilistic models have established a new state-of-the-art in generative imaging [6, 7]. They leverage a learned prior distribution of natural scenes. This way, diffusion models can effectively propose image details that are difficult to resolve solely from sparse noisy data. Diffusion models can then provide a way to handle non-linear inverse problems inherent to photon counting [10, 12].
2 THEORETICAL BACKGROUND
2.1 Langevin Dynamics
Consider a true physical object $\bm{x}$. It is randomly sampled from nature, with a natural Probability Density Function (PDF) denoted $p(\bm{x})$. We do not have hold of this object. We need to digitally generate an object (ideally, a natural one), by a process. A generated object is denoted $\bm{x}_{0}$. Generation seeks to maximize the probability $p(\bm{x}_{0})$.
In the Langevin dynamics algorithm, maximization of $p(\bm{x}_{0})$ is sought by a finite number of iterations, $K$. The process is initialized by an arbitrary unnatural object denoted $\bm{x}_{K}$. Then, let $k\in[K,\dots,1]$ be an iteration count-down index. A generated object at iteration $k$ is $\bm{x}_{k}$. Let $\bm{I}$ be the identity matrix and $\bm{\eta}_{k}\sim\mathcal{N}(\bm{0},\bm{I})$ be zero-mean, white Gaussian noise at iteration $t$. Then, generation by Langevin dynamics follows
$\bm{x}_{k-1}=\bm{x}_{k}+\alpha_{k}\nabla_{\bm{x}_{k}}\log p(\bm{x}_{k})+\sqrt{2\alpha_{k}}\bm{\eta}_{k}\;.$ (1)
Here $\alpha_{k}$ is a small non-negative value that decays with $k$, such that $\alpha_{k\to 1}=0$. Intuitively, Eq. (1) performs gradient ascent of the log-probability, gradually increasing the probability of $\bm{x}_{k}$. The added random $\bm{\eta}_{k}$ helps avoiding degeneration to a local maximum.
A generalization of this process solves inverse problems, as common in imaging. Let object $\bm{x}$ be imaged. The corresponding measured image data is noisy and denoted $\bm{y}$. Consider the posterior PDF of $\bm{x}$, given $\bm{y}$, which is denoted $p(\bm{x}|\bm{y})$. Now, solving an inverse problem is a generative process, which seeks to maximize $p(\bm{x}|\bm{y})$. In analogy to Eq. (1), the process is
$\bm{x}_{k-1}=\bm{x}_{k}+\alpha_{k}\nabla_{\bm{x}_{k}}\log p(\bm{x}_{k}|\bm{y})+\sqrt{2\alpha_{k}}\bm{\eta}_{k}\;.$ (2)
Following Bayes rule,
$\nabla_{\bm{x}_{k}}\log p(\bm{x}_{k}|\bm{y})=\nabla_{\bm{x}_{k}}\log p(\bm{x}_{k})+\nabla_{\bm{x}_{k}}\log p(\bm{y}|\bm{x}_{k})\;.$ (3)
From Eqs. (2,3),
$\bm{x}_{k-1}=\bm{x}_{k}+\alpha_{k}\nabla_{\bm{x}_{k}}\log p(\bm{x}_{k})+\alpha_{k}\nabla_{\bm{x}_{k}}\log p(\bm{y}|\bm{x}_{k})+\sqrt{2\alpha_{k}}\bm{\eta}_{k}\;.$ (4)
[…]
2.2 Approximating a Score Function
The score function of a PDF is $\nabla_{\bm{x}}\log p(\bm{x})$. Equations (1,4) require computation of the score function at each iteration. Unfortunately, we do not have access to $\nabla_{\bm{x}}\log p(\bm{x})$. To address this limitation, Ref. [8] introduces an estimator, known as the score network.
The score network is a deep neural network (DNN), having parameters $\bm{\phi}$. The input to the score network is an object $\bm{x}$. The network outputs a function $\bm{s}_{\bm{\phi}}(\bm{x})$. The DNN is trained so that $\bm{s}_{\bm{\phi}}(\bm{x})$ approximates the score function $\nabla_{\bm{x}}\log p(\bm{x})$. Training is supervised, using true objects $\bm{x}$, which are natural samples of $p(\bm{x})$. Training minimizes a loss function $C$. A desired loss may be of the form $\|\bm{s}_{\bm{\phi}}(\bm{x})-\nabla_{\bm{x}}\log p(\bm{x})\|_{2}^{2}$. In practice, Ref. [8] proposes a different loss $C$, whose minimization is equivalent, i.e., yielding the same $\bm{\phi}$ as the minimizer of $\|\bm{s}_{\bm{\phi}}(\bm{x})-\nabla_{\bm{x}}\log p(\bm{x})\|_{2}^{2}$.
Estimation using a score network has limitations. First, the score network $\bm{s}_{\bm{\phi}}(\bm{x})$ is not scalable to high dimensional objects, due to a costly computation [8],[9] of the loss function $C$. Second, $\bm{s}_{\bm{\phi}}(\bm{x})$ is unreliable for objects having a low probability $p(\bm{x})$. There are two reasons for this. (i) Hardly any training samples exist of objects having a low $p(\bm{x})$, thus under-constraining the DNN. (ii) In typical PDFs, the gradient $\nabla_{\bm{x}}\log p(\bm{x})$ is very low, at and around $\bm{x}$ for which $p(\bm{x})$ is low. Consequently, Eqs. (1,4) do not evolve properly. To address these limitations, Ref. [8] introduces annealed Langevin dynamics (ALD), described next.
2.3 Annealed Langevin Dynamics
[…]
A way to blur a PDF is to add a random vector (noise) $\bm{n}$ to a true object $\bm{x}$. Then, the true-object PDF is convolved with the PDF of the noise. Hence define an annealed object
$\tilde{\bm{x}}_{k}=\bm{x}+\bm{n}_{k}$ (5)
at iteration $k$, where $\bm{n}_{k}\sim\mathcal{N}(\bm{0},\sigma_{k}^{2}\bm{I})$. Generative iterations then evolve $\tilde{\bm{x}}_{k}$. Initially, being far from a natural object, the variance $\sigma_{K}^{2}$ is high, meaning that $\tilde{\bm{x}}_{K}$ is a very noisy scene. The blurred PDF $p_{\tilde{\bm{x}}}(\tilde{\bm{x}}_{K})$ is nevertheless effective at $\tilde{\bm{x}}_{K}$, leading to iterative refining of the scene. This refinement is akin to reversing a noise process. Hence, as iteration indices count-down from $K$ to 1, $\sigma_{k}^{2}$ decays. This process is ALD. Rather than Eqs. (1,4), it correspondingly uses
$\tilde{\bm{x}}_{k-1}=\tilde{\bm{x}}_{k}+\alpha_{k}\nabla_{\tilde{\bm{x}}_{k}}\log p_{\tilde{\bm{x}}}(\tilde{\bm{x}}_{k})+\sqrt{2\alpha_{k}}\bm{\eta}_{k},$ (6)
$\begin{split}\tilde{\bm{x}}_{k-1}=\tilde{\bm{x}}_{k}&+\alpha_{k}\nabla_{\tilde{\bm{x}}_{k}}\log p_{\tilde{\bm{x}}}(\tilde{\bm{x}}_{k})\\ &+\alpha_{k}\nabla_{\tilde{\bm{x}}_{k}}\log p(\bm{y}|\tilde{\bm{x}}_{k})+\sqrt{2\alpha_{k}}\bm{\eta}_{k}\;.\end{split}$ (7)
The parameter $\alpha_{k}$ is set per $k$ in accordance with $\sigma_{k}$. The true $\sigma_{k}$ is unknown, since the true $\bm{x}$ in Eq. (5) is unknown. Thus, an assumed step-wise function $\sigma_{k}$ is assumed, This function is part of the hyper-parameter set of the algorithm.
In analogy to Sec. 2.2, the score function $\nabla_{\bm{x}}\log p_{\tilde{\bm{x}}}(\tilde{\bm{x}}_{k})$ is approximated by the output $\bm{s}_{\bm{\theta}}(\tilde{\bm{x}}_{k})$ of a trained DNN. The DNN parameters are $\bm{\theta}$. However, the training cost $\tilde{C}$ is analytically tractable, being computationally more affordable for high-dimensional objects.
2.4 Likelihood in Non-Linear Problems
Eq. (7) provides a theoretical framework for inverse problems. However, calculating the likelihood gradient $\nabla_{\tilde{\bm{x}}_{k}}\log p(\bm{y}|\tilde{\bm{x}}_{k})$ may be a significant challenge for non-linear forward models. Recall that $\bm{y}=\mathcal{F}(\bm{x})$ is based on a true object. Now, consider how given a true object $\bm{x}$ and an annealed, noisy object $\tilde{\bm{x}}_{k}$, the PDFs $p(\bm{y}|\bm{x})$ and $p(\bm{y}|\tilde{\bm{x}}_{k})$ relate. They can often be related in a closed form in linear inverse problems. In non-linear cases, the dependency of $\bm{y}$ on the noisy state $\tilde{\bm{x}}_{k}$ is intractable, complicating computation of the gradient.
This matter is addressed by diffusion posterior sampling (DPS), proposed by Chung et al. [10]. Ref. [10] approximates the likelihood gradient using an estimate of the clean (not annealed) object, $\hat{\bm{x}}_{0}$. Using the ALD score-based formulation [8], this is equivalent to (see Ref. [11])
$\hat{\bm{x}}_{0}(\tilde{\bm{x}}_{k})\approx\tilde{\bm{x}}_{k}+\sigma_{k}^{2}\bm{s}_{\bm{\theta}}(\tilde{\bm{x}}_{k})\;.$ (8)
Then, the score approximation $\bm{s}_{\bm{\theta}}$ can be used to estimate $\nabla_{\tilde{\bm{x}}_{k}}\log p(\bm{y}|\tilde{\bm{x}}_{k})$ using the following connection,
$\nabla_{\tilde{\bm{x}}_{k}}\log p(\bm{y}|\tilde{\bm{x}}_{k})\approx\nabla_{\tilde{\bm{x}}_{k}}\log p[\bm{y}|\hat{\bm{x}}_{0}(\tilde{\bm{x}}_{k})]\;.$ (9)
This connection is closely related to Tweedie’s formula as presented in [10]. Taking the gradient $\nabla_{\tilde{\bm{x}}_{k}}$ of $\hat{\bm{x}}_{0}(\tilde{\bm{x}}_{k})$ is yielded via back-propagation through the DNN. Contrary to $\nabla_{\tilde{\bm{x}}_{k}}\log p(\bm{y}|\tilde{\bm{x}}_{k})$, the function $\nabla_{\hat{\bm{x}}_{0}}\log p(\bm{y}|\hat{\bm{x}}_{0})$ is easier to work with. The function $p(\bm{y}|\hat{\bm{x}}_{0})$ derives from the forward model $\mathcal{F}$. This formulation allows us to compute gradients through the non-linear forward model $\mathcal{F}$, using standard back-propagation.
3 Likelihood of Sequence of Events
3.1 Photon Detection Statistics
In this section we present the Poisson probability mass function (PMF) and Erlang PDF and cumulative distribution function (CDF), and then model photon detection events using both. Afterwards, we show how to relate the models. This relation is later used in Secs. 3.5, 3.6, 3.7 to prove the suggested likelihood is a valid probability distribution, by normalization to 1 with integration.
Notations (10)
$N\quad$ $\text{Number of observed events}.$
$\lambda\quad$ $\text{Photon detection event rate}.$
$\hat{\lambda}\quad$ $\text{Arbitrary rate for a Poisson process}.$
$\hat{t}\quad$ $\text{Arbitrary Erlang waiting time}.$
$t^{\prime}\quad$ $\text{Erlang integration variable}.$
$t_{N}\quad$ $\text{Random time variable of the $N$-th event}.$
$T\quad$ $\text{Exposure time}.$
$\tau_{d}\quad$ $\text{Dead time}.$
Figure 1: Visualizations of the Poisson and Erlang distributions in photon counting. [Top] The Poisson PMF for varying expected counts $\lambda T$, showing the probability of observing exactly NN events. [Middle] The Erlang PDF for varying NN, representing the distribution of detection time for the $N$-th event. [Bottom] The duality relationship showing that the Poisson probability (shaded region) is the difference between two Erlang CDFs.
The Poisson PMF (Fig. 1, Top) models the probability of observing exactly $N$ independent events occurring in an interval $\hat{t}$, given a known expected rate $\hat{\lambda}$:
$p(N)=\frac{(\hat{\lambda}\hat{t})^{N}\exp[-\hat{\lambda}\hat{t}]}{N!}.$ (11)
Here $N\in\{0,1,2,\dots\}$ represents the number of observed events [14].
The Erlang PDF (Fig. 1, Middle) describes the distribution of the waiting time $\hat{t}$ until the $N$-th event occurs in a Poisson process with rate $\hat{\lambda}$. It is a special case of the Gamma distribution, where the shape parameter is an integer NN[15, Ch. 5]:
$f_{\text{Erlang}}(\hat{t};N)=\frac{\hat{\lambda}^{N}\hat{t}^{N-1}\exp[-\hat{\lambda}\hat{t}]}{(N-1)!},\quad\hat{t}\geq 0.$ (12)
In a sensor unaffected by dead time $\tau_{d}$, a photo-detection event is modeled as a Poisson process with a constant rate parameter $\lambda$ (photons/second). Let $N$ denote the number of events within a fixed exposure time $T$. The probability of exactly $N$ events follows Eq. (11):
$p_{\text{Poisson}}(N)=\frac{(\lambda T)^{N}\exp[-\lambda T]}{N!},\quad N\in\{0,1,2,\dots\}\;.$ (13)
Alternatively, consider the continuous random variable $t_{N}$, representing the time of the $N$-th event. The PDF of $t_{N}$ follows Eq. (12):
$f_{\text{Erlang}}(t_{N};N)=\frac{\lambda^{N}t_{N}^{N-1}\exp[-\lambda t_{N}]}{(N-1)!},\quad t_{N}\geq 0\;.$ (14)
These two formulations are related. From Ref. [15, Sec. 5.3], having exactly $N$ events by time $T$ is equivalent to having the $N$-th event before time $T$, while the $(N+1)$-th event occurs after time $T$. Therefore, the Poisson PMF can be expressed as the difference between two Erlang CDFs (Fig. 1, Bottom):
$\begin{split}p_{\text{Poisson}}&(N)\\[10.0pt] &=\displaystyle\int_{0}^{T}f_{\text{Erlang}}(t^{\prime};N)\,dt^{\prime}-\int_{0}^{T}f_{\text{Erlang}}(t^{\prime};N+1)\,dt^{\prime}\\[10.0pt] &=F_{\text{Erlang}}(T;N)-F_{\text{Erlang}}(T;N+1)\;,\end{split}$ (15)
where
$F_{\text{Erlang}}(T;N)=1-\sum_{i=0}^{N-1}\frac{(\lambda T)^{i}\exp[-\lambda T]}{i!}\;.$ (16)
To verify the relation, Eq. (16) can be substituted back into Eq. (15):
$\begin{split}&\left(1-\sum_{i=0}^{N-1}\frac{(\lambda T)^{i}\exp[-\lambda T]}{i!}\right)-\left(1-\sum_{i=0}^{N}\frac{(\lambda T)^{i}\exp[-\lambda T]}{i!}\right)\\[10.0pt] &=\sum_{i=0}^{N}\frac{(\lambda T)^{i}\exp[-\lambda T]}{i!}-\sum_{i=0}^{N-1}\frac{(\lambda T)^{i}\exp[-\lambda T]}{i!}\;.\end{split}$ (17)
The first sum runs up to $N$, while the second sum runs up to N1N-1. So, all terms cancel out, except for the term where $i=N$. This leaves:
$\frac{(\lambda T)^{N}\exp[-\lambda T]}{N!},$ (18)
which is precisely the Poisson PMF definition from Eq. (13).
3.2 Inter-arrival Events Statistics
In this section we discuss the distribution of the time between consecutive events. It is used in Sec. 3.4 as part of the suggested forward model.
Notations (19)
$\Delta t\quad$ $\text{Time between two consecutive events}.$
$\lambda\quad$ $\text{Photon detection event rate}.$
The time interval $\Delta t$ between consecutive events in a Poisson process follows an exponential distribution [15, Ch. 5]. The probability for having no event during $\Delta t$ is
$p({\rm no~event})=\exp[-\lambda\Delta t]\;.$ (20)
The CDF for having an event during $\Delta t$ is then
$F_{\text{exp}}(\Delta t)=1-p({\rm no~event})=1-\exp[-\lambda\Delta t].$ (21)
The derivative of Eq. (21) is the exponential PDF,
$f_{\text{exp}}(\Delta t)=\lambda\exp[-\lambda\Delta t]\;.$ (22)
3.3 Photon Detection with Dead Time
This section describes dead time and how it affects photon detection event statistics. We show how this affects the exponential probability formula of Sec. 3.2.
Notations (23)
$\Delta t\quad$ $\text{Time difference between two consecutive events}.$
$\tau_{d}\quad$ $\text{Dead time}.$
$\lambda\quad$ $\text{Photon detection event rate if }\tau_{d}=0.$
Figure 2: The shifted exponential distribution representing the inter-detection time $\Delta t$ in the presence of dead time. The standard exponential distribution (gray dashed line, $\tau_{d}=0$) is shifted by the dead time period $\tau_{d}$, creating a region $0,\tau_{d})$ where the probability of detection is zero.
In a SPAD pixel, each detection triggers a dead time $\tau_{d}$ during which the sensor is insensitive to incoming photons. The time interval between the end of a dead time window and the next detection event is exponentially distributed, with the same rate $\lambda$ as the incident Poisson process [13].
From Ref. [13], the inter-detection time $\Delta t$ follows a shifted exponential PDF as seen in Fig. 2:
$\begin{split}f_{\text{exp}}^{\text{shifted}}&(\Delta t)\\[10.0pt] &=\begin{dcases}\lambda\exp\left[-\lambda(\Delta t-\tau_{d})\right]&\Delta t\geq\tau_{d}\\[10.0pt] 0&\Delta t<\tau_{d}\;.\end{dcases}\end{split}$ (24)
3.4 Forward Model
In this section we define a forward model. We describe a probabilistic model for SPAD signals, based on the measured sequence of events, and a given photon flux. This model leads to a likelihood function.
Notations (25)
$N\quad$ $\text{Total number of detected photon events}.$
$q\quad$ $\text{Quantum efficiency}.$
$T\quad$ $\text{Exposure time}.$
$\Phi\quad$ $\text{Photon flux}.$
$\{t_{i}\}_{i=1}^{N}\quad$ $\text{Set of $N$ event detection times}.$
$\Psi\quad$ $\text{Definition for $\prod_{i=2}^{N}\lambda\exp\left[-\lambda(t_{i}-t_{i-1}-\tau_{d})\right]$}.$
$\lambda\quad$ $\text{Photon detection event rate if }\tau_{d}=0.$
$\tau_{d}\quad$ $\text{Dead time}.$
A forward model is a mathematical operator that maps an underlying object to observed measurements, accounting for the physics of the sensing system. Here we want a model that relates photon flux to a measured sequence of photon detection events.
Let $\Phi$ denote the incident photon flux and $q\in[0,1]$ be the quantum efficiency of the pixel. If there is no dead-time, the effective photon detection rate is
$\lambda=q\Phi\;.$ (26)
Let $T$ be the exposure time of a SPAD pixel. The total number of photons detected by the pixel during this interval is denoted by the random variable $N$. The detection event times are denoted by a set of random variables $\{t_{i}\}_{i=1}^{N}$, where $0\leq t_{1}<\dots<t_{N}\leq T$. Due to the dead time, the time intervals satisfy
$t_{i}-t_{i-1}\geq\tau_{d}\quad\forall i\in\{2,\dots,N\}\;.$ (27)
The PDF for detecting the first photon at time $t_{1}$ is described by Eq. (22)
$\lambda\exp[-\lambda t_{1}]\;.$ (28)
The PDF of each of the consequent event times is derived by Eq. (24). So, their joint PDF is
$\Psi=\prod_{i=2}^{N}\lambda\exp\left[-\lambda(t_{i}-t_{i-1}-\tau_{d})\right]\;.$ (29)
The probability that no detection events happened in the remaining exposure time is derived by Eq. (20)
$\exp\left[-\lambda(T-t_{N}-\tau_{d})\right]\;.$ (30)
The probability no detection event happenes at all during exposure time $T$ is also derived by Eq. (20)
$\exp\left[-\lambda T\right]\;.$ (31)
Eqs. (28, 29, 30) form the PDF, or likelihood, to have a set of detection events described by $(N,\{t_{i}\}_{i=1}^{N})$, given $\lambda$ and $N>0$:
$\begin{split}p(N,&\{t_{i}\}_{i=1}^{N}|\lambda)\\[10.0pt] &=\lambda\exp[-\lambda t_{1}]\Psi\exp[-\lambda(T-t_{N}-\tau_{d})]\;.\end{split}$ (32)
In the instance of no detection events $N=0$, the likelihood is described by Eq. (31)
$p(N=0|\lambda)=\exp\left[-\lambda T\right]\;.$ (33)
To verify that Eqs. (32, 33) constitute a valid PDF, they must satisfy two conditions: non-negativity and normalization.
The first condition,
$\begin{split}p(N,\{t_{i}\}_{i=1}^{N}|\lambda)&\geq 0,\quad N\in\{1,2,\dots\}\\ p(N=0|\lambda)&\geq 0\end{split}$ (34)
is satisfied by definition. Since the rate parameter $\lambda$ is positive and the exponential function is strictly positive for all real arguments, the likelihood function, constructed as a product of these terms, is strictly positive.
The second condition requires that a sum over all $N$, of the respective integrals over the domain of event times equal unity:
$\sum_{N=0}^{\infty}\left[\int_{t_{N}=0}^{T}\dots\int_{t_{2}=0}^{t_{3}}\int_{t_{1}=0}^{t_{2}}p(N,\{t_{i}\}_{i=1}^{N}|\lambda)\,dt_{1}\dots dt_{N}\right]=1\;.$ (35)
The proof is split into two cases according to the arrival time $t_{N}$ of the last measured event.
3.5 Case I
In this section we formulate the likelihood in the case where the time left after the $N$-th event is longer than $\tau_{d}$. Then we present the result in the Erlang form along Eq. (15)
Notations (36)
$\mathcal{A}_{N}\quad$ $\text{Definition for $\int_{0}^{u_{N}}\dots\int_{0}^{u_{2}}1\ du_{1}\dots du_{N-1}$}.$
$N\quad$ $\text{Total number of detected photon events}.$
$\{t_{i}\}_{i=1}^{N}\quad$ $\text{Set of $N$ event detection times}.$
$T\quad$ $\text{Exposure time}.$
$\{u_{i}\}_{u=1}^{N}\quad$ $\text{Change of variables use for $\{t_{i}\}_{t=1}^{N}$}.$
$\Omega\quad$ $\text{Definition for }\prod_{i=1}^{N}\lambda\exp\left[-\lambda(u_{i}-u_{i-1})\right].$
$\tau_{d}\quad$ $\text{Dead time}.$
$\lambda\quad$ $\text{Photon detection event rate if }\tau_{d}=0.$
Figure 3: Case I: The NN-th event occurs early enough ($t_{N}\leq T-\tau_{d}$) that the dead time ends before the exposure time $T$, leaving a window where further detection is possible.
Assuming at least one detection event $N>0$. As shown in Fig. 3, after the last measured event $t_{N}$ and its corresponding dead-time period $\tau_{d}$, there is still time left until the exposure time $T$ is exhausted. Therefore,
$0<t_{N}\leq T-\tau_{d}\;.$ (37)
Define
$\begin{split}&u_{0}=0\\ &u_{i}=t_{i}+(1-i)\tau_{d}\quad\quad\text{for $i\in\{1,2,\dots\,N\}$}\;.\end{split}$ (38)
Then,
[…]
The variables $\{u_{i}\}_{i=2}^{N}$ satisfy
$\begin{split}u_{i}-u_{i-1}&=[t_{i}-(i-1)\tau_{d}]-[t_{i-1}-(i-2)\tau_{d}]\\[10.0pt] &=t_{i}-t_{i-1}-\tau_{d}\;.\end{split}$ (40)
Assigning back to Eq. (29), define
$\begin{split}\Omega&=\lambda\exp[-\lambda u_{1}]\prod_{i=2}^{N}\lambda\exp\left[-\lambda(u_{i}-u_{i-1})\right]\\ &=\lambda\exp[-\lambda(u_{1}-u_{0})]\prod_{i=2}^{N}\lambda\exp\left[-\lambda(u_{i}-u_{i-1})\right]\\ &=\prod_{i=1}^{N}\lambda\exp\left[-\lambda(u_{i}-u_{i-1})\right]\end{split}$ (41)
Following Eq. (32),
$\begin{split}p(N,&\{u_{i}\}_{i=1}^{N}|\lambda)\\[10.0pt] &=\Omega\exp[-\lambda\{T-(u_{N}+[N-1]\tau_{d})-\tau_{d}\}]\\[10.0pt] &=\lambda^{N}\exp[-\lambda u_{N}]\exp[-\lambda(T-N\tau_{d}-u_{N})]\\[10.0pt] &=\lambda^{N}\exp[-\lambda(T-N\tau_{d})]\;.\end{split}$ (42)
Integrating over $\{u_{i}\}_{i=1}^{N}$ to calculate $p(N)$ under the integral bounds $0<u_{1}<u_{2}<\dots<u_{N}<(T-N\tau_{d})$,
$\begin{split}p(&N)=\int_{u_{N}=0}^{T-\tau_{d}}\dots\int_{u_{1}=0}^{u_{2}}\lambda^{N}\exp[-\lambda(T-N\tau_{d})]du_{1}\dots du_{N}\\[10.0pt] &=\lambda^{N}\exp[-\lambda(T-N\tau_{d})]\int_{0}^{T-\tau_{d}}\dots\int_{0}^{u_{2}}1\ du_{1}\dots du_{N}\;.\end{split}$ (43)
Denote
$\mathcal{A}_{N}=\int_{u_{N-1}=0}^{u_{N}}\dots\int_{u_{1}=0}^{u_{2}}1\ du_{1}\dots du_{N-1}\;.$ (44)
Integrating Eq. (44) for different values of $N$ gives:
$\begin{split}\\ \mathcal{A}_{1}&=1\\[10.0pt] \mathcal{A}_{2}&=\displaystyle\int_{0}^{u_{2}}1\,du_{1}=u_{2}\\[10.0pt] \mathcal{A}_{3}&=\displaystyle\int_{0}^{u_{3}}u_{2}\,du_{2}=\frac{u_{3}^{2}}{2}\\ &\vdots\\ \mathcal{A}_{N}&=\displaystyle\int_{0}^{u_{N}}\frac{u_{N-1}^{N-2}}{(N-2)!}\,du_{N-1}=\frac{u_{N}^{N-1}}{(N-1)!}\;.\end{split}$ (45)
The final integral over $\mathcal{A}_{N}$ follows
$\begin{split}\displaystyle\int_{0}^{T-N\tau_{d}}\mathcal{A}_{N}\,du_{N}&=\displaystyle\int_{0}^{T-N\tau_{d}}\frac{u_{N}^{N-1}}{(N-1)!}\,du_{N}\\[10.0pt] &=\frac{(T-N\tau_{d})^{N}}{N!}\;.\end{split}$ (46)
Substituting Eq. (46) back to Eq. (43),
$\begin{split}p(N)&=\lambda^{N}\exp[-\lambda(T-N\tau_{d})]\frac{(T-N\tau_{d})^{N}}{N!}\\[10.0pt] &=\frac{[\lambda(T-N\tau_{d})]^{N}\exp[-\lambda(T-N\tau_{d})]}{N!}\;.\end{split}$ (47)
We note that for $N=0$, Eq. (47) coincides with Eq. (33)
$\frac{[\lambda(T-0\cdot\tau_{d})]^{0}\exp[-\lambda(T-0\cdot\tau_{d})]}{0!}=\exp{[-\lambda T]}\;.$ (48)
Therefore, Eq. (47) is true for $N\in\{0,1,2,\dots\}$. Using Eq. (15),
$p(N)=F_{\text{Erlang}}(T-N\tau_{d};N)-F_{\text{Erlang}}(T-N\tau_{d};N+1)\;.$ (49)
3.6 Case II
In this section we formulate the likelihood in the case where the time left after the $N$-th event is shorter than $\tau_{d}$. We then present the result in the Erlang form along Eq. (15)
Notations (50)
$\mathcal{A}_{N}\quad$ $\text{Definition for $\int_{0}^{u_{N}}\dots\int_{0}^{u_{2}}1\ du_{1}\dots du_{N-1}$}.$
$N\quad$ $\text{Total number of detected photon events}.$
$\{t_{i}\}_{i=1}^{N}\quad$ $\text{Set of $N$ event detection times}.$
$\{u_{i}\}_{u=1}^{N}\quad$ $\text{Change of variables use for $\{t_{i}\}_{t=1}^{N}$}.$
$T\quad$ $\text{Exposure time}.$
$\tau_{d}\quad$ $\text{Dead time}.$
$\lambda\quad$ $\text{Photon detection event rate if }\tau_{d}=0.$
$\Psi\quad$ $\text{Definition for $\prod_{i=2}^{N}\lambda\exp\left[-\lambda(t_{i}-t_{i-1}-\tau_{d})\right]$}.$
Figure 4: Case II: The NN-th event occurs late ($t_{N}>T-\tau_{d}$), meaning the dead time window extends beyond $T$, preventing any further detection.
Assuming $N>0$. After the final measured event $t_{N}$, the time remaining before exhaustion of the exposure time $T$ is shorter than $\tau_{d}$. So,
$T-\tau_{d}<t_{N}\leq T\;.$ (51)
There can thus be no feasibility for an additional photon detection, as visualized in Fig. 4. The probability of having no detection event after $t_{N}$ is 1. Eq. (32) simplifies to
$p(N,\{t_{i}\}_{i=1}^{N}|\lambda)\\ =\lambda\exp[-\lambda t_{1}]\Psi\;.$ (52)
Using change of variables from Eq. (38),
$\begin{split}p(N,\{u_{i}\}_{i=1}^{N}|\lambda)&=\prod_{i=1}^{N}\lambda\exp[-\lambda(u_{i}-u_{i-1})]\\[10.0pt] &=\lambda^{N}\exp[-\lambda\sum_{i=1}^{N}(u_{i}-u_{i-1})]\\[10.0pt] &=\lambda^{N}\exp[-\lambda u_{N}]\;.\end{split}$ (53)
Under change of variables from Eq. (38) , Eq. (51) becomes
$T-N\tau_{d}<u_{N}\leq T-(N-1)\tau_{d}\;.$ (54)
Integrating over $\{u_{i}\}_{i=1}^{N}$ to calculate $p(N)$ under the integral bounds $0<u_{1}<u_{2}<\dots<u_{N}<[T-(N-1)\tau_{d}]$, with the same methodology as in Eqs. (44, 45),
$\begin{split}p(N)&=\int_{u_{N}=T-N\tau_{d}}^{T-(N-1)\tau_{d}}\dots\int_{u_{1}=0}^{u_{2}}\lambda^{N}\exp[-\lambda u_{N}]du_{1}\dots du_{N}\\[10.0pt] &=\int_{T-N\tau_{d}}^{T-(N-1)\tau_{d}}\lambda^{N}\exp[-\lambda u_{N}]\mathcal{A}_{N}du_{N}\\[10.0pt] &=\int_{T-N\tau_{d}}^{T-(N-1)\tau_{d}}\lambda^{N}\exp[-\lambda u_{N}]\frac{u_{N}^{N-1}}{(N-1)!}du_{N}\;.\end{split}$ (55)
The integrand in Eq. (55) is the Erlang PDF shown in Eq. (14). Hence,
$p(N)=F_{\text{Erlang}}[T-(N-1)\tau_{d};N]-F_{\text{Erlang}}[T-N\tau_{d};N]\;.$ (56)
In the instance of no detection events $N=0$, there is no dead time $\tau_{d}$ affecting the model. Therefore, the situation described by $\tt{case\ II}$ cannot occur and
p(N=0)=0.p(N=0)=0\;. (57)
We note that $F_{\text{Erlang}}[T;N=0]$ and $F_{\text{Erlang}}[T+\tau_{d};N=0]$ are the probabilities to detect at least 0 events by times $T$ and $T+\tau_{d}$ respectively:
$\begin{split}F_{\text{Erlang}}&(T;N=0)\\ &=1-\cancelto{0}{\sum_{i=0}^{0-1}}\frac{(\lambda T)^{i}\exp[-\lambda T]}{i!}=1\\[10.0pt] F_{\text{Erlang}}&(T+\tau_{d};N=0)\\ &=1-\cancelto{0}{\sum_{i=0}^{0-1}}\frac{[\lambda(T+\tau_{d})]^{i}\exp[-\lambda(T+\tau_{d})]}{i!}=1\;.\end{split}$ (58)
Assigning $N=0$ in Eq. (56) and using Eq. (58) gives
$\begin{split}p(N=0)&=F_{\text{Erlang}}(T+\tau_{d};N=0)-F_{\text{Erlang}}(T;N=0)\\ &=1-1=0\;,\end{split}$ (59)
which coincides with Eq. (57). Therefore, Eq. (56) is true for $N\in\{0,1,2,\dots\}$.
3.7 Combined Case
In sections 3.5, 3.6, the likelihood function is derived for two popssible classes of detection time $t_{N}$, in relation to the exposure time $T$ and dead time $\tau_{d}$. In this section, both likelihoods are combined to represent the overall likelihood. This is and then integrated to 1, to prove it is a valid PMF.
Notations (60)
$M\quad$ $\text{Maximum number of detection events possible}.$
$N\quad$ $\text{Total number of detected photon events}.$
$S_{N}\quad$ $\text{Definition for $F_{\text{Erlang}}[T-(N-1)\tau_{d};N]$}.$
$T\quad$ $\text{Exposure time}.$
$\tau_{d}\quad$ $\text{Dead time}.$
The combined probability distribution function $p(N)$ can be calculated:
$\begin{split}&p(N)=p_{\tt{CaseI}}(N)+p_{\tt{CaseII}}(N)\\[10.0pt] &=\{F_{\text{Erlang}}[T-N\tau_{d};N]-F_{\text{Erlang}}[T-N\tau_{d};N+1]\}\\[10.0pt] &+\{F_{\text{Erlang}}[T-(N-1)\tau_{d};N]-F_{\text{Erlang}}[T-N\tau_{d};N]\}\\[10.0pt] &=F_{\text{Erlang}}[T-(N-1)\tau_{d};N]\\ &\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -F_{\text{Erlang}}[T-N\tau_{d};N+1]\;.\end{split}$ (61)
The maximum theoretical number $M$ of detection events by exposure time $T$ occurs when detections occur immediately after a dead time has passed,
$M\triangleq\left\lfloor\frac{T}{\tau_{d}}\right\rfloor+1\;.$ (62)
Let $S_{N}$ be defined as
$S_{N}\triangleq F_{\text{Erlang}}[T-(N-1)\tau_{d};N]\;.$ (63)
By the definition of the Erlang CDF,
[…]
Using Eqs. (61, 62, 63, 64) then proves Eq. (35)
$\begin{split}\sum_{N=0}^{M}{p(N)}&=(S_{0}-S_{1})+(S_{1}-S_{2})+\dots+(S_{M}-S_{M+1})\\[10.0pt] &=S_{0}-S_{M+1}\\[10.0pt] &=1-0\\[10.0pt] &=1\end{split}$ (65)
3.8 Likelihood Score
In this section we rely on sections 3.5, 3.6, and define the likelihood term. Then, the term of the likelihood score is determined by applying a logarithm, and deriving a gradient with respect to $\lambda$.
Notations (66)
$N\quad$ $\text{Total number of detected photon events}.$
$\{t_{i}\}_{i=1}^{N}\quad$ $\text{Set of $N$ event detection times}.$
$T\quad$ $\text{Exposure time}.$
$\lambda\quad$ $\text{Photon detection event rate if~}\tau_{d}=0.$
$\Psi\quad$ $\text{Definition for $\prod_{i=2}^{N}\lambda\exp\left[-\lambda(t_{i}-t_{i-1}-\tau_{d})\right]$}.$
$\tau_{d}\quad$ $\text{Dead time}.$
The likelihood of the measurements $(N,\{t_{i}\}_{i=1}^{N})$ for a given $\lambda$ is
$\begin{split}p(N,&\{t_{i}\}_{i=1}^{N}|\lambda)\\[10.0pt] &=\begin{cases}\displaystyle\lambda\exp[-\lambda t_{1}]\Psi\exp\left[-\lambda(T-t_{N}-\tau_{d})\right]&\tt{Case\ I}\\[10.0pt] \displaystyle\lambda\exp[-\lambda t_{1}]\Psi&\tt{Case\ II}\;,\end{cases}\end{split}$ (67)
where
${\tt Case\ I:}\ \ \ 0<t_{N}\leq T-\tau_{d}$
${\tt Case\ II:}\ \ \ T-\tau_{d}<t_{N}\leq T$
. The following summation term
$\begin{split}\sum_{i=2}^{N}(t_{i}-t_{i-1})\;,\end{split}$ (68)
simplifies to
$\begin{split}t_{N}-t_{1}\;.\end{split}$ (69)
Using Eq. (69) on Eq. (29),
$\begin{split}\displaystyle\prod_{i=2}^{N}\lambda&\exp\left[-\lambda(t_{i}-t_{i-1}-\tau_{d})\right]\\[10.0pt] &=\lambda^{N-1}\exp\left[-\lambda\sum_{i=2}^{N}(t_{i}-t_{i-1}-\tau_{d})\right]\\[10.0pt] \displaystyle&=\lambda^{N-1}\exp\left[-\lambda\{t_{N}-t_{1}-(N-1)\tau_{d}\right)\}]\;.\end{split}$ (70)
Substituting Eq. (70) in Eq. (67):
[…]
The log likelihood is
$\begin{split}\log p(N,\{t_{i}\}_{i=1}^{N}|\lambda)&=\begin{cases}\displaystyle\log\{\lambda^{N}\exp[-\lambda(T-N\tau_{d})]\}&\tt{I}\\[10.0pt] \displaystyle\log\{\lambda^{N}\exp[-\lambda(t_{N}-N\tau_{d})]\}&\tt{II}\end{cases}\\[10.0pt] &=\begin{cases}\displaystyle N\log\{\lambda\}-\lambda(T-N\tau_{d})&\tt{I}\\[10.0pt] \displaystyle N\log\{\lambda\}-\lambda(t_{N}-N\tau_{d})&\tt{II}\;.\end{cases}\end{split}$ (72)
The gradient of the log likelihood in Eq. (72) with respect to $\lambda$ is
$\begin{split}\nabla_{\lambda}\log p(N,\{t_{i}\}_{i=1}^{N}|\lambda)&=\begin{cases}\displaystyle\frac{N}{\lambda}-N\left(\frac{T}{N}-\tau_{d}\right)&\tt{I}\\[10.0pt] \displaystyle\frac{N}{\lambda}-N\left(\frac{t_{N}}{N}-\tau_{d}\right)&\tt{II}\end{cases}\end{split}$ (73)
4 Domain Adaptation
In this section we bridge the domain difference between the estimation $\hat{\bm{x}}_{0}$ described in Sec. 2 and the input of the forward model defined in Sec. 3.4. Then, we relate the likelihood score function to the input to the diffusion model.
Notations (74)
$a\quad$ $\text{Domain adaptation scaling factor}.$
$b\quad$ $\text{Domain adaptation bias}.$
$\bm{N}\quad$ A vector of the number of measured detection
[…]
$s_{\bm{\theta}}\quad$ $\text{Diffusion model score function}.$
[…]
$\bm{\lambda}\quad$ A vector of photon flux per pixel,
$\text{across the FOV}.$
A discrepancy exists between the operational domains of the diffusion model and the SPAD signal. The score network $s_{\bm{\theta}}$ and its resulting estimate $\hat{\bm{x}}_{0}$ operates in a normalized range of [1,1][-1,1]. In contrast, the SPAD forward model $\mathcal{F}$ depends on a physical photon flux $\bm{\lambda}$ (photons per second) of Eq. (26), which is non-negative. Let $b$ act as a bias that ensures values of $\hat{\bm{x}}_{0}$ are none-negative. Let $a$ be a scaling factor that converts grayscale units to physical photon flux. We define an affine transformation in the form of
$\bm{\lambda}=a(\hat{\bm{x}}_{0}+b)\;.$ (75)
Following Eq. (75), we compute $\nabla_{\bm{\lambda}}\log p\!\left(\bm{N},\{t_{i}\}_{i=1}^{\bm{N}}\mid\bm{\lambda}\right)$. We then multiply this gradient by $\rho_{k}$, which both normalizes it and incorporates the required chain-rule factor induced by the domain transformation from $\hat{\bm{x}}_{0}$ to $\bm{\lambda}$.
5 Reconstruction Algorithm
In this section we provide a pseudo-code for the reconstruction of an object using a sequence of photon detection events. Events are measured by a SPAD. The pseudo-code also uses a diffusion model trained to generate data from the distribution of the object.
Algorithm 1 Diffusion Posterior Sampling (DPS) Algorithm
1:Measurements $(\bm{N},\{t_{i}\}_{i=1}^{\bm{N}})$, trained score network $\bm{s}_{\bm{\theta}}$, forward model parameters, schedules $\{\alpha_{k},\beta_{k},\sigma_{k},\rho_{k}\}_{k=0}^{K}$.
2:Reconstructed image $\hat{\bm{x}}_{0}$.
3:Initialize $\bm{x}_{K}\sim\mathcal{N}(\bm{0},\bm{I})$
4:for $k=K-1\textbf{ to }0$ do
5:  1. Score Estimation:
6:  $\hat{\bm{s}}\leftarrow\bm{s}_{\bm{\theta}}(\bm{x}_{k},k)$
7:  2. Tweedie’s Estimation (Clean Image):
8:  $\hat{\bm{x}}_{0}\leftarrow\frac{1}{\sqrt{\alpha_{k}}}\left[\bm{x}_{k}+(1-\alpha_{k})\hat{\bm{s}}\right$]
9:  3. Sample Noise:
10:  $\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})$
11:  4. Ancestral Step (Reverse Diffusion):
12:  $\bm{x}^{\prime}_{k-1}\leftarrow\frac{\sqrt{\alpha_{k}}(1-\alpha_{k-1})}{1-\alpha_{k}}\bm{x}_{k}+\frac{\sqrt{\alpha_{k-1}}\beta_{k}}{1-\alpha_{k}}\hat{\bm{x}}_{0}+\sigma_{k}\bm{z}$
13:  5. Domain Adaptation:
14:  $\bm{\lambda}\leftarrow a(\hat{\bm{x}}_{0}+b)$
15:  6. Guidance Step (Likelihood Gradient):
16:  $\bm{x}_{k-1}\leftarrow\bm{x}^{\prime}_{k-1}-\rho_{k}\nabla_{\bm{\lambda}}\log p(\bm{N},\{t_{i}\}_{i=1}^{\bm{N}}|{\bm{\lambda}})$
17:end for
18:return $\bm{\lambda}$
6 Tests
6.1 Simulation of the Forward Model
Figure 5: Simulation of photon detection events for different lux values. For visualization, a simple reconstruction method is used.
In this section we detail a simulation, which generates measurements of photon detection events, based on grayscale images. To simulate consecutive detection events, we use the method detailed in Ref. [16]. The method takes grayscale images with values normalized to the range $[0,1]$ and maps them to photon flux. The parameters we used for the simulation are: Pixel pitch $5[\mu\text{m}]$, $q=0.9$), $\tau_{d}=50[ns]$), fill factor 1, wavelength $555[nm]$, luminous efficiency $683{[\text{lm/W}]}$, dark count rate $100[\text{Hz}]$, after-pulsing probability ($100[\text{ns}]$), $\sigma_{\rm jitter}=200[\text{ps}]$, exposure time $1[\text{ms}]$. The reference lux differs across tests. In Fig. 5 an example is shown of a simulation of photon detection events corresponding to an image with different lux levels, but the same exposure time. A simple reconstruction method from Ref. [13] is used for visualization.
6.2 Results
In this section we show reconstruction results. The tests use a diffusion model from Ref. [10], pre-trained on the FFHQ dataset [17]. We use two approaches.
Approach A
Here all pixels have a uniform exposure time $T$. We measure the sequence of detection events for each pixel. In this method, each pixel may record a different number of events. Then, we reconstruct the image using the algorithm in Sec. 5. Example of reconstructions for different simulated values of photon flux appear in Figs. 6,7.
Approach B
First, a desired number of detection events $N_{\text{det}}$ is chosen. We simulated each SPAD sensor pixel with the same parameter, but an unlimited exposure time. Exactly $N_{\text{det}}$ detections are registered per pixel. Then, $\bm{\lambda}$ is reconstructed using the algorithm in Sec. 5. Example of reconstructions based on $N_{\text{det}}=100$ appear in Fig. 8. Reconstructions based on $N_{\text{det}}=1$ appear in Fig. 9.
Figure 6: Approach A: Three example images. They were used to simulate photon detection sequences using 400[mlux] as the reference lux simulation input. Metrics used are PSNR (better $\uparrow$) and LPIPS (better $\downarrow$). From top to bottom: Reconstructed images, reference images, simple reconstruction using Ref. [13], histogram of detected events.
Figure 7: Approach A: Three example images. They were used to simulate photon detection sequences using 40[mlux] as the reference lux simulation input. Metrics used are PSNR (better $\uparrow$) and LPIPS (better $\downarrow$). From top to bottom: Reconstructed images, reference images, simple reconstruction using Ref. [13], histogram of detected events.
Figure 8: Approach B: Three example images. They were used to simulate photon detection sequences of 100 events per pixel. Metrics used are PSNR (better $\uparrow$) and LPIPS (better $\downarrow$). From top to bottom: Reconstructed images, reference images.
Figure 9: Approach B: Three example images. They were used to simulate the detection time of just one (the first) photon per pixel. Reconstruction was performed using the single (and first) time of arrival only per pixel. Metrics used are PSNR (better $\uparrow$) and LPIPS (better $\downarrow$). From top to bottom: Reconstructed images, reference images.